Honda Car Purchase and Geek Squad - SMy Pages-South MAT2058 Statistics SUO2 RoseJefferson 1 1/15/19 10:30 Test: Week 5 Final Exam This Question: 15 pts 8 of 20 (6 complete) ▼ This Test: 275 pts p 漫Qu Below are 36 sorted ages of an acting award winner. Find P10 using the method presented in the textbook 18 41 18 42 19 57 61 65 65 Enter your answer in the answer box. บาน2019 Type here to search
Answer #1

18

18

19

20

22

27

31

31

39

40

41

42

46

46

48

49

52

52

53

54

55

56

56

57

62

62

62

62

63

63

65

65

66

74

80

step 1. Arrange the data in ascending order.

Total 35 no.

P(10)= 0.1 * 35 = 3.5th no.

P(10)= (19+20)/2 = 19.5

P(10) = 19.5